# Josephson anomalous vortices

### v1

@@ Abstract
We show that vortices with circulating current, related with odd-frequency triplet pairing, appear in Josephson junctions where the barrier is a weak ferromagnet with strong spin-orbit coupling. By both symmetry analysis and microscopic methods we show that there is an additional term — a rotary invariant — in the superconducting free energy which allows for magnetoelectric effects even when the previously considered Lifshitz invariant vanishes. We show that the size, shape, and position of these vortices can be controlled by manipulating Rashba spin-orbit coupling in the weak link, via gates, and we suggest that these vortices could be detected via scanning magnetometry techniques. We also show that the transverse triplet components of the superconducting correlations can form a texture.
@@
Introduction— Vortices in superconductors have remained a vibrant area of research since their discovery in the 1950s. In the original work by Abrikosov [1], vortices appear in type II superconductors threaded by a flux quantum as a result of the superconducting phase winding by $2\pi$ around a singularity. This produces a signature circulating current with a normal core. Superconducting heterostructures allow for a rich variety of other kinds of vortices. For example, in Josephson junctions coreless, elongated vortices can arise due to kinks or solitons in the superconducting phase [2, 3]. Vortices with cores can also appear in normal metals in proximity to a superconductor [4], and also in junctions due to destructive interference of the superconducting correlations [5]. While in these cases, an external magnetic field is required to generate vortices, other forms of magnetism can also generate vortices. For example, evidence has been presented for vortices in Fe(Se, Te) due to the internal magnetization [6, 7], and on superconducting surfaces with strong spin-orbit coupling (SOC) and magnetic adatoms [8]. In the latter case, spin-triplet superconducting correlations are created by the ferromagnetic proximity effect, and modulated by spin-orbit coupling. In analogy to superfluid He-3 [9, 10, 11], it might be possible to generate also more exotic vortex matter with such triplet correlations.
Figure 1: (a) Sketch of a superconductor-ferromagnetic metal-superconductor device studied here. There is strong spin-orbit coupling in the ferromagnetic layer. (b) Circulating current density in the weak link of (a) due to the rotary invariant. Parameters: $(L_{x},L_{y},J,\alpha)=(\xi_{0},\xi_{0},\Delta_{0},2.35/\xi_{0})$.
@@
Here we introduce the Josephson anomalous vortex (JAV). These appear as a circulating current density (Fig. missing 1(b)) in two-dimensional junctions where the Josephson weak link is ferromagnetic and has strong SOC (Fig. missing 1(a)), as long as no current is applied. JAVs only appear when the weak link, or gate-defined regions in the weak link, are close to a coherence length in length and width. The current density is at a maximum at the edges of the vortex, and vanishes at the center. Complementary to these current vortices, we also present two kinds of superconducting textures: (i) a texture in the phase of the $f_{z}$ component, and (ii) a texture formed from the amplitudes of the $f_{x}$ and $f_{y}$ components of the superconducting correlations. The latter are analogous to a magnetic texture.
@@
In this Letter we study a previously unexplored rotary invariant in the free energy. We show that this invariant generates circulating currents, even when the standard Lifshitz invariant [12] vanishes. We show that JAVs are not fine-tuned so that they can appear over a range of spin-orbit and exchange energies. We argue that, due to the magnetic field produced by swirling currents, JAVs could be detected by scanning magnetometry techniques. We also suggest that these vortices may be manipulated by Rashba “islands” within the weak link via back gates.
@@
Throughout this Letter we use the singlet superconducting gap $\Delta_{0}$ of the electrodes as the natural unit of energy, and the superconducting coherence length $\xi_{0}$ as the unit of length. We restore the physical units when predicting amplitudes of measurable quantities. In numerical calculations we fix the temperature to $T=0.1\Delta_{0}/k_{\rm B}$. We use the conventions that Greek subscripts $\mu,\nu$ run over $(0,1,2,3)$, while Latin subscripts run over $(1,2,3)$. Repeated indices imply summation.
@@
Magnetoelectric effects without a Lifshitz invariant— Over the past decades, it has been shown experimentally and theoretically that SOC allows for coupling between the magnetic and electric properties of a material [12]. Such coupling is expressed in phenomena such as the superconducting and Josephson diode effects [13], helical phases [14], and the Edelstein and inverse Edelstein effects [15, 16, 17]. For weak superconductivity many magnetoelectric effects can be described by the addition of a Lifshitz invariant in the Ginzburg-Landau free energy [12]. In this Letter we investigate a rotary invariant in the free energy which also gives rise to magnetoelectric effects, even when there is no Lifshitz invariant. This term was reported in [18], but its significance was not discussed.
@@
$\mathcal{F}=\,$ $a|\psi|^{2}+\frac{b}{2}|\psi|^{4}+\frac{1}{2}D_{jk}\Pi_{j}\psi(\Pi_{k}\psi)^{*}+\frac{|{\bm{B}}|^{2}}{2\mu_{0}}$
$+d_{k}[\psi^{*}\Pi_{k}\psi+\psi(\Pi_{k}\psi)^{*}]$
$+ie_{jk}\Pi_{j}\psi(\Pi_{k}\psi)^{*}.$ (1)
@@
Here $\psi$ is the complex order parameter of the superconductor, $a$ and $b$ are the usual Ginzburg-Landau constants, $D_{jk}$ the superfluid stiffness symmetric tensor, and ${\bm{B}}$ the magnetic field. For compactness we define the momentum $\Pi_{j}=-i\hbar\hat{\nabla}_{j}=-i\hbar\partial_{j}-2eA_{j}$, with $A_{j}$ the electromagnetic vector potential. In addition to the conventional Ginzburg-Landau terms, symmetries allow several additional terms [19]: the Lifshitz invariant vector $d_{k}$, and the rotary invariant antisymmetric tensor $e_{jk}$ which is of interest here.
@@
The presence of $d_{k}$ and $e_{jk}$ requires breaking of time-reversal symmetry. The vector $d_{k}$ also requires intrinsic breaking of inversion symmetry. The $e_{jk}$ term is a total derivative, and requires breaking of inversion symmetry by extrinsic inhomogeneity, for example system boundaries (see Appendix A for details). One mechanism for breaking those symmetries comes from the combination of SOC and exchange field ${\bm{h}}$, which we consider here. In the case of linear in momentum SOC, with $H_{\rm SO}=\mathcal{A}_{j}^{a}k_{j}\sigma^{a}/(2m)$, we can define the SU(2) fields $\mathcal{A}_{j}=\mathcal{A}_{j}^{a}\sigma^{a}$ [20, 21] and the corresponding field strength tensor $F_{ij}=-i[\mathcal{A}_{i},\mathcal{A}_{j}]$, valid for position independent SU(2) fields. Using these definitions, we find that (see Sec. S2 in [22]) generally
$d_{i}=i\kappa{\rm Tr}\left\{F_{ij}[\mathcal{A}_{j},h_{k}\sigma_{k}]\right\},\quad e_{ij}=\kappa{\rm Tr}[F_{ij}h_{k}\sigma_{k}],$ (2)
@@
where $\kappa=-\frac{7\zeta(3)\nu_{\rm F}\eta}{16\pi^{2}k_{B}^{2}T_{c}^{2}}$, with $\nu_{\rm F}$ the density of states per spin projection at the Fermi level, $\eta=D\ell^{2}/(k_{\rm F}\ell)$, $D=v_{\rm F}\ell/2$ the diffusion constant, $v_{\rm F}$ the Fermi velocity, $k_{\rm F}$ the Fermi wavevector, $T_{c}$ is the critical temperature, and $\ell$ the mean free path. $\zeta(n)$ is the Zeta function. Thus $d_{k}$ is cubic and $e_{ij}$ quadratic in the strength of SOC. Moreover, these forms imply that $d_{k}$ is non-zero only when the exchange field and $F_{ij}$ have perpendicular components, whereas $e_{ij}$ requires the presence of their parallel components. We can hence study the implications of one without the other.
@@
In the setup shown in Fig. missing 1(a), with an exchange field ${\bm{h}}=J\hat{{\bm{z}}}$ and Rashba SOC due to an electric field also in the $z$ direction, the SOC Hamiltonian is $H_{\rm SO}=\alpha(\hat{{\bm{z}}}\times{\bm{k}})\cdot{\bm{\sigma}}$ and the corresponding field-strength tensor is $F_{xy}=-F_{yx}=2\alpha^{2}\sigma_{z}$. Hence, $d_{k}=0$ and $e_{xy}=-e_{yx}=4\kappa\alpha^{2}J$. Integration by parts of this rotary term results in a boundary term $\propto{}n_{j}e_{jk}[\psi\partial_{k}\psi^{*}-\mathrm{c.c.}]$ that generates circulating currents perpendicular to the boundary normal $n_{j}$.
We derive the rotary invariant from microscopics in the diffusive regime, based on the methods described in Ref. 23, 24 (see Sec. S1 in [22] for details). While Eq. (missing) 1 is valid for a bulk superconducting system, for heterostructures such as the one sketched in Fig. missing 1(a) the corresponding rotary invariant of the microscopic theory is
@@
Here $\zeta=\pi\nu_{\rm F}\hbar\eta\alpha^{2}$; in all numerical results we assume $\hbar\eta=0.01\Delta_{0}\xi_{0}^{4}$. As this is a proximity system we write the free energy in terms of superconducting correlations ${\bm{f}}=f_{s}\sigma_{0}+f_{i}\sigma_{i}$, and due to the ferromagnetic proximity effect we require that the correlations have both singlet $f_{s}$ and triplet ${\bm{f}}_{t}=(f_{x},f_{y},f_{z})$ components [25]. We assume that the ferromagnet is weak, so that the quasiclassical approximation is valid (i.e., the density of states is the same for both up and down spins). The rotary term $\mathcal{F}_{\rm rot}$ appears as long as there is both SOC and an exchange field, even when there is no Lifshitz invariant. It generates circulating current densities,
@@
We numerically solve the full nonlinear problem, including singlet-triplet conversion and finite-size effects, using the methods described in Ref. 18. In Fig. missing 1(b) we plot the current density for a junction of $L_{x}\times L_{y}=\xi_{0}\times\xi_{0}$, at zero phase bias. This yields a representative current vortex, distinguished by the strong current density at the boundary (as opposed to the center, as in an Abrikosov vortex). For large junctions we find some swirling current close to the superconducting terminals [22], similar to what has been reported in Ref. 26 for in-plane exchange fields. As long as $\alpha^{2}\xi_{0}^{2}\gg\varphi$, the rotary invariant, and thus the circulating current, dominate over the usual Josephson current. For $\varphi\gtrsim\alpha^{2}\xi_{0}^{2}$ the rotary invariant perturbs the Josephson current (see Sec. S4 in [22]).
Figure 2: Magnetic field strength along the $\hat{z}$ direction at the center of the junction due to circulating currents for various junction sizes, over a range of exchange energies $J$ and spin-orbit coupling strengths $\alpha$. Dashed lines indicate the transitions (T) between antivortex (AV) and vortex (V) phases. Inset: representative plot of an antivortex. Parameters for inset: $(L_{x},L_{y},J,\alpha)=(\xi_{0},\xi_{0},\Delta_{0},1/(2\xi_{0}))$.
Magnetic field due to a vortex— Using the Biot-Savart law, we plot in Fig. missing 2 the magnetic field at the centre of the vortex, for four different junction sizes at zero phase bias, over a range of exchange energies $J$ and SOC strengths $\alpha$. The magnetic field scale in the plot is $B_{0}=\mu_{0}\sigma_{\rm D}\Delta_{0}t/(4e\xi_{0})=1$ mT×$\times(t/\xi_{0})[\sigma_{\rm D}/(1.6\cdot 10^{7}\text{ S/m})][\Delta_{0}/(200\text{ $\mu$eV})]$, expressed in terms of relatively typical values for the Drude conductivity $\sigma_{\rm D}=\nu_{\rm F}e^{2}D$ in disordered metals, and a typical $\Delta_{0}$ for Al. We also include the thickness $t\ll\xi_{0}$ of the junction. We do not compute this magnetic field self-consistently (i.e., by feeding back into the electromagnetic vector potential). Due to competition between the rotary invariant and other terms in the free energy, current vortices can only appear above some threshold $\alpha,J$, depending, e.g., on boundary conditions, temperature, and field strength. We show in Appendix B, for example, that from the Ginzburg-Landau free energy vortices are stable when the rotary invariant exceeds the superfluid weight. Numerically, we find a vortex- antivortex transition as a function of $\alpha$ and $J$ (we plot a representative antivortex in the inset of Fig. missing 2). The antivortex phase arises due to the boundary conditions, with the exact dependence on $\alpha$ and $J$ varying with system size. Note that in the antivortex phase there is no net current transport across the junction and no associated magnetic field.
@@
${\bm{J}}_{\rm V}=J_{0}\left(\frac{-y}{\xi_{0}}\hat{{\bm{x}}}+\frac{x}{\xi_{0}}\hat{{\bm{y}}}\right),$ ${\bm{J}}_{\rm AV}=J_{0}\left(\frac{y}{\xi_{0}}\hat{{\bm{x}}}+\frac{x}{\xi_{0}}\hat{{\bm{y}}}\right),$ (5)
with $J_{0}=e\zeta\Delta_{0}/(\hbar\xi_{0})$. It follows then that for an antivortex the circulation is $\Gamma_{\rm AV}=0$, but for a vortex the circulation is $\Gamma_{\rm V}=2J_{0}A$, with $A$ the area of the vortex. By computing $\Gamma$ over a square patch in the center of the weak link in the vortex phase, we have that $J_{\rm max}\approx J_{0}$. Hence in the numerics $\Gamma\approx 2A$ indicates a vortex, $\Gamma\approx 0$ an antivortex, and $\Gamma\approx A$ the transitional phase. Due to finite size effects and differences between the square geometry as compared to a disc, the numerical vorticity is not quantized and we see small variations compared to the analytic result.
In polar coordinates, the approximate current density for a vortex is ${\bm{J}}_{\rm V}=J_{0}\rho/\xi_{0}\hat{{\bm{\phi}}}$, with $\rho$ the radial coordinate and $\phi$ the angular coordinate. The magnetic field at the center of the vortex ($\rho=0$) is then

### v2

@@ Abstract
We show that vortices with circulating current, related with odd-frequency triplet pairing, appear in Josephson junctions where the barrier is a weak ferromagnet with strong spin-orbit coupling. By symmetry analysis we show that there is an additional term — a rotary invariant — in the superconducting free energy which allows for magnetoelectric effects even when the previously considered Lifshitz invariant vanishes. Using a microscopic model based on a modified Usadel equation incorporating those effects, we show that the size, shape, and position of these vortices can be controlled by manipulating Rashba spin-orbit coupling in the weak link, via gates, and we suggest that these vortices could be detected via scanning magnetometry techniques. We also show that the transverse triplet components of the pairing amplitudes can form a texture.
@@
Introduction— Vortices in superconductors have remained a vibrant area of research since their discovery in the 1950s. In the original work by Abrikosov [1], vortices appear in type II superconductors threaded by a flux quantum as a result of the superconducting phase winding by $2\pi$ around a singularity. This produces a signature circulating current with a normal core. Superconducting heterostructures allow for a rich variety of other kinds of vortices. For example, in Josephson junctions coreless, elongated vortices can arise due to solitons in the superconducting phase [2]. Vortices with cores can also appear in normal metals in proximity to a superconductor [3], and also in junctions due to destructive interference of the pairing amplitudes [4]. Swirling supercurrents and vortices can appear in structures with strong spin-orbit coupling (SOC), for example when there are magnetic impurities [5] or when an in-plane exchange field is applied [6, 7]. In these cases, spin-triplet pairing amplitudes are created by the exchange coupling, and modulated by spin-orbit coupling. In analogy to superfluid He-3 [8, 9, 10], it might be possible to generate also more exotic vortex matter in superconducting systems with such triplet pairing amplitudes.
@@
Here we introduce the Josephson anomalous vortex (JAV). These appear as a circulating current density (Fig.  1(b)) in two-dimensional junctions where the Josephson weak link is ferromagnetic and has strong SOC (Fig.  1(a)), as long as no current is applied. JAVs only appear when the weak link, or gate-defined regions in the weak link, are close to a coherence length in length and width. The current density is at a maximum at the edges of the vortex, and vanishes at the center. Correspondingly, at the center of a vortex the triplet part of the pairing amplitudes vanishes, leaving only the singlet part; these vortices do not have a core. These amplitudes are textured in a manner similar to a magnetic vortex.
Figure 1: (a) Sketch of a superconductor-ferromagnetic metal-superconductor device studied here. There is strong spin-orbit coupling in the ferromagnetic layer. (b) Circulating current density in the weak link of (a) due to the rotary invariant, computed numerically using the microscopic model. Here no current is applied across the junction. Parameters: $(L_{x},L_{y},J,\alpha)=(\xi_{0},\xi_{0},\Delta_{0},2.35/\xi_{0})$.
@@
In this Letter we study a previously unexplored rotary invariant in the free energy. The rotary invariant is related to a boundary term, and so in the absence of an applied current generates circulating currents. Circulating currents due to the rotary invariant arise even when the standard Lifshitz invariant [11] vanishes. We show that JAVs are not fine-tuned so that they can appear over a range of spin-orbit and exchange energies. We argue that, due to the magnetic field produced by swirling currents, JAVs could be detected by scanning magnetometry techniques. We also suggest that these vortices may be manipulated by Rashba “islands” within the weak link via back gates.
@@
Throughout this Letter we use the singlet superconducting gap $\Delta_{0}$ of the proximity induced superconductivity in the ferromagnet as the natural unit of energy, and the superconducting coherence length $\xi_{0}$ as the unit of length. We restore the physical units when predicting amplitudes of measurable quantities. In numerical calculations we fix the temperature to $T=0.1\Delta_{0}/k_{\rm B}$. We use the conventions that Greek subscripts $\mu,\nu$ run over $(0,1,2,3)$, while Latin subscripts run over $(1,2,3)$. Repeated indices imply summation.
@@
Magnetoelectric effects without a Lifshitz invariant— Over the past decades, it has been shown experimentally and theoretically that SOC allows for coupling between the magnetic and electric properties of a material [11]. Such coupling is expressed in phenomena such as the superconducting and Josephson diode effects [12], helical phases [13], and the Edelstein and inverse Edelstein effects [14, 15, 16]. For weak superconductivity many magnetoelectric effects can be described by the addition of a Lifshitz invariant in the Ginzburg-Landau free energy [11]. In this Letter we show first from symmetry that magnetoelectric effects can give rise to boundary invariants in addition to the bulk Lifshitz invariant in the most generic free energy. The most interesting is an invariant that we term the rotary invariant (reported earlier in Ref. [17] but with its significance not discussed). Such rotary invariants can appear in a wide range of superconducting systems when there is no Lifshitz invariant, as long as the relevant symmetries are satisfied, and generally they describe vortex-like excitations appearing in inhomogeneous systems or systems with boundaries. With insight from this general symmetry analysis, we apply a concrete microscopic model (using quasiclassical methods [18]) describing disordered superconductors in the presence of spin-orbit coupling, derive the microscopic expressions for the invariants, and study the vortices in parameter regimes much beyond the Ginzburg-Landau theory.
@@
$\mathcal{F}=\mathcal{F}_{\rm S}+\mathcal{F}_{\rm L}+\mathcal{F}_{\rm R}.$ (1)
Here $\mathcal{F}_{\rm S}$ is the usual Ginzburg-Landau free energy [19],
$\mathcal{F}_{\rm S}=a|\psi|^{2}+\frac{b}{2}|\psi|^{4}+\frac{1}{2}D_{jk}\Pi_{j}\psi(\Pi_{k}\psi)^{*}+\frac{|{\bm{B}}|^{2}}{2\mu_{0}},$ (2)
@@
with $\psi$ the complex order parameter of the superconductor, $a$ and $b$ the Ginzburg-Landau constants, $D_{jk}$ the superfluid stiffness symmetric tensor, and $\bm{B}$} the magnetic field. For compactness we define the momentum $\Pi_{j}=-i\hbar\hat{\nabla}_{j}=-i\hbar\partial_{j}-2eA_{j}$, with $A_{j}$ the electromagnetic vector potential. If time-reversal symmetry is broken, and inversion symmetry intrinsically broken, we can add $d_{k}$, the Lifshitz invariant vector, via the second term $\mathcal{F}_{\rm L}$ [20], with
$\mathcal{F}_{\rm L}=d_{k}[\psi^{*}\Pi_{k}\psi+\psi(\Pi_{k}\psi)^{*}].$ (3)
If time-reversal symmetry is broken and the inversion symmetry is broken by extrinsic inhomogeneity (e.g., via system boundaries, see Appendix A) we can add the rotary invariant $e_{jk}$, an antisymmetric tensor, via the third term $\mathcal{F}_{\rm R}$, with
$\mathcal{F}_{\rm R}=ie_{jk}\Pi_{j}\psi(\Pi_{k}\psi)^{*}.$ (4)
@@
One mechanism for breaking time-reversal and inversion symmetry comes from the combination of SOC and exchange field $\bm{h}$}, which we consider here. In the case of linear in momentum SOC, with $H_{\rm SO}=\mathcal{A}_{j}^{a}k_{j}\sigma^{a}/(2m)$, we can define the SU(2) fields $\mathcal{A}_{j}=\mathcal{A}_{j}^{a}\sigma^{a}$ [21, 22] and the corresponding field strength tensor $F_{ij}=-i[\mathcal{A}_{i},\mathcal{A}_{j}]$, valid for position independent SU(2) fields. Using these definitions, we find that (see Sec. II in the supplement [23]) generally
$d_{k}=i\kappa\,{\rm Tr}\left\{F_{kj}[\mathcal{A}_{j},h_{i}\sigma_{i}]\right\},\quad e_{jk}=\kappa\,{\rm Tr}[F_{jk}h_{i}\sigma_{i}],$ (5)
@@
where $\kappa=-\frac{7\zeta(3)\nu_{\rm F}\eta}{16\pi^{2}k_{B}^{2}T_{c}^{2}}$, with $\nu_{\rm F}$ the density of states per spin projection at the Fermi level, $\eta=D\ell^{2}/(k_{\rm F}\ell)$, $D=v_{\rm F}\ell/2$ the diffusion constant, $v_{\rm F}$ the Fermi velocity, $k_{\rm F}$ the Fermi wavevector, $T_{c}$ is the critical temperature, and $\ell$ the mean free path. $\zeta(n)$ is the Zeta function. Thus $d_{k}$ is cubic and $e_{jk}$ quadratic in the strength of SOC (here we assume weak SOC; in contrast, when $\alpha\,k_{\rm F}\gg\Delta_{0}$ the Lifshitz invariant $d_{k}$ is linear in $\alpha$[13]). Moreover, these forms imply that $d_{k}$ is non-zero only when the exchange field and $F_{ij}$ have perpendicular components, whereas $e_{jk}$ requires the presence of their parallel components. We can hence study the implications of one without the other.
@@
To illustrate this point, we focus on the junction geometry shown in Fig.  1(a). Here we have an exchange field $^{\bm{h}}=J\hat{{\bm{z}}}$, and Rashba SOC due to an electric field also in the $z$ direction. The SOC Hamiltonian is $H_{\rm SO}=\alpha(\hat{{\bm{z}}}\times{\bm{k}})\cdot{\bm{\sigma}}$ and the corresponding field-strength tensor is $F_{xy}=-F_{yx}=2\alpha^{2}\sigma_{z}$. Hence, the Lifshitz invariant vanishes: $d_{k}=0$, and the rotary invariant is finite: $e_{xy}=-e_{yx}=4\kappa\alpha^{2}J$. Integration by parts of this rotary term results in a boundary term $\propto{}n_{j}e_{jk}[\psi\partial_{k}\psi^{*}-\mathrm{c.c.}]$ that generates circulating currents perpendicular to the boundary normal $n_{j}$.
To investigate the rotary invariant outside of the Ginzburg-Landau regime, for the rest of this Letter we study a microscopic model for the heterostructure sketched in Fig.  1(a), using the methods described in Refs. 18, 24, 25 (see Sec. I in the supplement [23] for details). From the preceding analysis we expect a finite rotary invariant (and the related magnetoelectric effects) for this geometry because the relevant symmetries are satisfied. We assume that the ferromagnet is weak, so that the quasiclassical approximation is valid (i.e., the density of states is the same for both up and down spins). Note that we assume that the superconducting electrodes proximity induce superconductivity into the ferromagnetic thin film region just below the electrodes. In general this gap $\Delta_{0}$ is smaller than the zero-temperature, zero-field gap of the electrodes. Spin-orbit coupling also increases the critical field, so we can study the effects of large exchange field, $J\sim\Delta_{0}$ [26].
Due to the exchange field, the order parameter (i.e., pairing amplitude or anomalous Green’s function) for this system is not scalar, but rather has a spin structure, ${\bm{f}}=f_{s}\sigma_{0}+f_{i}\sigma_{i}$ [27]. Here $f_{s}$ is the singlet part and $f_{i}={\bm{f}}_{t}=(f_{x},f_{y},f_{z})$ the triplet part of the anomalous Green’s function. As expected from the Ginzburg-Landau analysis, this microscopic model also has a rotary invariant. For this model, the rotary invariant involves the $f_{s}$ and $f_{z}$ components of the pairing amplitude (see Sec. I in the supplement [23]),
@@
Here $\zeta=\pi\nu_{\rm F}\hbar\eta\alpha^{2}$; in all numerical results we assume $\hbar\eta=0.01\Delta_{0}\xi_{0}^{4}$. Note that this term scales as $\sim 1/E_{\rm F}$ and so breaks the quasiclassical symmetry [28, 24] (that is, the combined time-reversal and electron-hole symmetries). The rotary term $\mathcal{F}_{\rm rot}$ appears as long as there is both SOC and an exchange field, even when there is no Lifshitz invariant. It generates circulating current densities,
@@
We emphasize that these circulating currents are distinct to those introduced in Ref. [29], which are due to an interfacial SOC generated surface Lifshitz invariant. While these surface currents can be similar effect here. This depends on the geometry: in Fig.  1(a), the FM/S interface normal is parallel to the exchange field, and the interfacial Lifshitz invariant vanishes. In contrast, in a lateral junction (where the ferromagnetic weak link is sandwiched between two superconductors on the left and right), both the interface and the bulk effects would matter.
Figure 2: (a) Texture formed from the anomalous Green’s functions $f_{x},f_{y}$, reminiscent of a magnetic texture. (b) Phase winding of the anomalous Green’s functions $f_{x}+if_{y}$, reminiscent of an Abrikosov vortex. (c) Phase texture of the superconducting correlation $f_{z}$. (d) Representative antivortex pattern in the current density. (e) Representative mixture between a current antivortex and a vortex patterns. (f) Representative vortex. Parameters for all panels: $(L_{x},L_{y},J)=(\xi_{0},\xi_{0},\Delta_{0})$; for (a-c): $\alpha=2/\xi_{0}$; for (d): $\alpha=1/\xi_{0}$; for (e): $\alpha=1.55/\xi_{0}$; for (f): $\alpha=2/\xi_{0}$;
@@
In Fig.  1(b) we plot the current density for a junction of $L_{x}\times L_{y}=(1\times 1)\xi_{0}^{2}$, at zero phase bias. This yields a representative current vortex, distinguished by the strong current density at the boundary (as opposed to the center, as in an Abrikosov vortex) and vanishing current in the center. We compute numerically current densities and pairing amplitudes for the microscopic model using the code described in Ref. 17. For large junctions we find some swirling current close to the superconducting terminals (see Sec. III in the supplement [23]), similar to what has been reported in Ref. 7 for in-plane exchange fields. As long as $\alpha^{2}\xi_{0}^{2}\gg\varphi$, the rotary invariant, and thus the circulating current, dominate over the usual Josephson current. For $\varphi\gtrsim\alpha^{2}\xi_{0}^{2}$ the rotary invariant perturbs the Josephson current (see Sec. IV in the supplement [23]).
Figure 3: (a-d) Total Poincaré index $\mathcal{P}$ for various junction sizes, over a range of exchange energies $J$ and spin-orbit coupling strengths $\alpha$. Antivortex patterns in the current correspond to $\mathcal{P}=-1$ and vortices to $\mathcal{P}=+1$, as indicated by the inset sketches. Junction size is indicated in the top right hand corner, $L_{x}\times L_{y}\xi_{0}^{2}$. (e-h) Magnetic field strength along the $\hat{z}$ direction at the centre of the junction, for the same parameter spaces. The Poincaré index and magnetic field strength are computed from the numerically calculated current densities.
Antivortex patterns and superconducting textures — Often current vortices in superconductors are accompanied by textures in the phase of the superconducting order parameter; in this case the anomalous Green’s function (pairing amplitude) $\bm{f}$}. The connection between current vortex patterns and pairing amplitudes in the case considered here is more complicated. Namely, for all finite exchange fields $J$ and Rashba spin-orbit coupling strengths $\alpha$ tested we find numerically that the triplet part of the pairing amplitudes is textured. In Fig.  2(a) we plot $f_{x},f_{y}$ as a vector field, forming a swirling pattern reminiscent of a magnetic texture. Equivalently, we plot in Fig.  2(b) the complex phase of the $f_{x}+if_{y}$ anomalous Green’s functions, featuring a phase winding reminiscent of an Abrikosov vortex. Finally, we plot in Fig.  2(c) the phase texture of the $f_{z}$ triplet. This texture does not completely wind, due to the line defect along $x=0$. All three textures presented are representative, with only small variations with $J,\alpha$. In all our simulations, the amplitudes scale like $|f_{s}|>|f_{z}|>|f_{x}|\sim|f_{y}|$. From Eq. () 7 it is clear that the $f_{z}$ texture generates current vortex patterns (Fig.  2(f)). In addition to vortices we also find antivortex patterns (Fig.  2(d)) [30] in the current density, and mixtures between the two (Fig.  2(e)). As with the vortices, the current vanishes in the center of an antivortex pattern. The appearance of the latter patterns is due to the influence of the $f_{x},f_{y}$ textures, and depends subtly on the small variations of ${\bm{f}}_{t}$ with $J,\alpha$.
@@
The most natural way to categorize singularities in vector fields (i.e., points in the weak link where the current density vanishes) is with the Poincaré index. This is the winding number [31]
$\mathcal{P}=\frac{1}{\pi}\oint_{C}d\theta,$ (8)
with $\theta$ the angle a vector field (such as the current density) makes relative to the $x$-axis, and $C$ a loop around a singularity. When there are multiple singularities, the total index is the sum of all $\mathcal{P}$ for each singularity. This quantity is a good topological invariant, because small variations in $\bm{J}$} do not change $\mathcal{P}$ [32]. However, because the current density does not vanish far from the singularity, we see a continuous transition between antivortex and vortex patterns, rather than a first-order transition.
To gain some intuition for this index, consider the current densities ${\bm{J}}_{\rm V}$ and ${\bm{J}}_{\rm AV}$, which approximate the vortex and antivortex patterns (at least not too close to the boundaries),
[…]
Here $J_{0}=e\zeta\Delta_{0}/(\hbar\xi_{0})$. By choosing the integration path $C$ to be a circle, a straightforward calculation shows that $\mathcal{P}=-1$ for antivortex and $\mathcal{P}=+1$ for vortex patterns — this is true regardless of the vorticity (this can also be seen by noting that the change in the angle, with fixed integration direction, does not depend on vorticity).
We plot in Fig.  3(a-d) the total index over a range of $J,\alpha$, for 4 different system sizes, $L_{x}\times L_{y}=\{0.1\times 0.1,0.5\times 0.5,1\times 1,5\times 5\}\xi_{0}^{2}$. For the systems smaller than $L_{x}\times L_{y}=(5\times 5)\xi_{0}^{2}$ there are only two phases, vortex and antivortex (as sketched in the inset). For the largest system tested, $(5\times 5)\xi_{0}^{2}$, the phase diagram is richer. For $\alpha>1.25/\xi_{0}$ there are multiple singularities, leading to a total index $\mathcal{P}_{\rm tot}>1$. This results in more complicated patterns in the current density that cannot be simply categorized as a vortex or an antivortex [23]. When there is a single singularity ($\alpha<1.25/\xi_{0}$) Eq. () 9 is valid and we find vortex and antivortex patterns in the current density. The smallest system tested, $(0.1\times 0.1)\xi_{0}^{2}$, does not in general support vortices, preferring antivortices. We find that vortices are preferred when the system size is close to $L_{x}=L_{y}=\xi_{0}$. Clearly system size is essential to realizing current vortices, as the current density is related to gradients of the anomalous Green’s functions $\bm{f}$}. For the same reason the appearance of antivortex rather than vortex patterns depends subtly on small variations in ${\bm{f}}_{t}$.
[…]